OpenStudy (anonymous):

Show Newton's Method to approximate the real zero of the function in the interval [0,2]. Give the answer to the nearest .001

OpenStudy (anonymous):

f(x)=x^5+x-1

OpenStudy (anonymous):

@yun2thejae

OpenStudy (anonymous):

Oh boy.

OpenStudy (anonymous):

give me a second.

OpenStudy (anonymous):

do you know the newton's method? I'm asking because i want to map out how much I have to explain

OpenStudy (anonymous):

yeah \[x _{1} - \frac{ f(x _{1}) }{ f'(x _{1} )}\] @yun2thejae

OpenStudy (anonymous):

k that is good. generally, you want to repeat newton's method a few times to get really close to the actual root. Choose a number for xsub1 so that you can find xsub2. Anyway, start with xsub1=1. Plug in the value and solve. and then repeat. so if xsub1=1 then xsub2= 1- (f(1)/f'(1))

OpenStudy (anonymous):

so choose numbers between [0,2] for \[x _{1}\]

OpenStudy (anonymous):

well you only need to choose one number. easiest number to pick is 1.

OpenStudy (anonymous):

yeah i chose 1 and i got .833

OpenStudy (anonymous):

when you find xsub2, use xsub2 and plug it in to find xsub3. et cetera et cetera.

OpenStudy (anonymous):

now do .833-f(.833)/f'(.833)

OpenStudy (anonymous):

or just keep pressing enter in ur calculator until it stops producing new numbers lol

OpenStudy (anonymous):

ok the number it stopped producing at was .7548776662

OpenStudy (anonymous):

do it once or twice more and use that as your answer. yeah. when I did this in class, we weren't allowed to use a graphing calculator at all during the semester

OpenStudy (anonymous):

so my answer is .755 :)

OpenStudy (anonymous):

but then for the other question for this part of my final \[f(x)=\sqrt[3]{x-1} \] my teacher put no answer, does that mean that this particular equation doesn't work with Newton's Method? :|

OpenStudy (anonymous):

@yun2thejae

OpenStudy (anonymous):

might mean that there are no real roots.

OpenStudy (anonymous):

realy? ok

OpenStudy (anonymous):

I'm not really. sure. that would be my first assumption.

OpenStudy (zarkon):

newtons method does not wok with the above equation even though there is clearly a root at x=1

OpenStudy (anonymous):

oh ok thx :)

OpenStudy (zarkon):